# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v3

@@ 3.2 Stage 2: Net Fragments
Here, the forward connections have the shape $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)},C^{(S1)},h,w}$ and the recurrent connections have the shape $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)},C^{(S2)},h,w}$, where $C^{(S2)}$ is the number of channels in $S2$. Therefore, the membrane potential $\mathcal{M}$ at time $t$ is increased by the learned amount $w_{j,i}$ from all connected and active neurons ($y_{j,t-1}=1$). Note that $\mathcal{F}_{0}(n_{i,0})$ only consists of neurons of $S1$ since all neurons in $S2$ are initialized to silent state ($y_{i,0}=0$).
Higher support, indicated by larger $a^{\mathcal{M}(S2)}_{i}$, suggests a more likely pattern, as neurons encoding frequent patterns gain reinforcement not only from feedforward connections but also from lateral connections. Before we discuss how the weight matrices $\boldsymbol{W}^{(F)},\boldsymbol{W}^{(L)}$ are initialized and trained, we describe how the membrane potential $\boldsymbol{a}^{\mathcal{M}(S2)}$ undergoes attenuation and is transformed into $\boldsymbol{y}^{(S2)}$.
@@
The membrane potential is processed by an activation function $\mathcal{A}\left(\boldsymbol{a}^{\mathcal{M}},\mathcal{E}\right)$. In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{M}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range $[0,1]$ (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active. Specifically, attenuation is implemented as:
@@
Alternative Neurons.
A single neuron that encodes a specific feature can be activated in multiple contexts. For example, a short straight-line segment may appear in various contexts – such as in different line drawings or digits. In our framework, each context is represented by a net, whereas individual features correspond to the activity of specific neurons.
To achieve context disentanglement, it is important to allocate sufficient representational capacity such that the same feature (e.g., a line segment) can be utilized across distinct contexts (e.g., different line drawings) without interference. This requires that a feature-representing neuron is capable of participating in multiple, context-specific net fragments. Without this capacity, a neuron would be forced to average over competing contextual associations, resulting in a non-specific or “blurry” representation.
Formally, a neuron representing a feature may be part of multiple net fragments: $n_{i}\in\mathcal{F}(n_{j})$ and $n_{i}\in\mathcal{F}(n_{k})$ where $n_{j}\neq n_{k}$). This implies that the neuron must form distinct lateral connections $w_{i,j}$ and $w_{i,k}$, which should not be concurrently active.
To address this requirement, we introduce “alternative neurons”. Each alternative version of a neuron is activated by the same feature from $S1$ but has a distinct set of lateral connections in $S2$ to encode different contexts. This is implemented by increasing the channel dimension $C^{(S2)}$ of $\boldsymbol{W}^{(L)}$ by a factor $\kappa$, i.e.,
$C^{(S2)}:=C^{(S2)}\cdot\kappa$ (19)
This expansion provides capacity for $\kappa$ alternative version of a neuron. Consequently, these alternative neurons receive identical bottom-up input from $S1$ while differing in the lateral support they receive in $S2$. A winner-take-all (WTA) competition mechanism ensures that each alternative neuron specializes in a distinct context, with only one being activated in any given instance.
@@
At initialization, neuron selection is random. However, after Hebbian learning updates, the alternative neurons progressively specialize, leading to distinct and structured fragment representations. Over time, the most correlated alternative neuron is chosen for each input location, ensuring robust feature specialization across different contexts.
Weight Initialization.
The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are initially copied into all alternative $S2$ neurons in the same position. This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to $1$ if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of alternative neurons. The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to $1$.
Learning of Weights.
Support within net fragments is learned via Hebbian plasticity (Hebb, 1949), so that frequently occurring patterns that lead to correlated activation patterns between neurons $n_{i}\sim n_{j}$, are “stored” in the system and receive more support:
Here, $\eta$ represents the learning rate and $\boldsymbol{W}^{(S2)}$ stands for both $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$. Note that the same convolutional kernel is applied at various spatial positions. The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects (across all spatial positions). If the resulting average is positive, the corresponding connection increases as it connects more simultaneously active neurons than neurons that fire disjointly. Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced. Learning is applied after processing a sample (after $T$ timesteps) to improve stability.
4 Experiments
To demonstrate the compositionality of net fragments into nets, we train the model on binary images containing straight-line patterns so that the CNA learns net fragments corresponding to straight-line sub-patterns (smaller straight-line segments). However, for evaluation (after training and without weight updates), we use much more complex images such as line drawings of objects, and demonstrate that these complex patterns can be expressed as compositions (nets) of straight-line net fragments. This flexible compositional behavior contrasts strongly with conventional associative memory that cannot model such out-of-distribution data (see Section 6.1). To highlight the differences to ANNs, we train and evaluate an autoencoder on the same dataset and compare the representations.
We also examine robustness by introducing noise. Specifically, we assess the efficacy of filtering out random Gaussian noise and reconstructing partially occluded objects. Further experimental details are provided in the appendix: We describe the training and testing datasets in more detail in Appendix D, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ is described in Appendix A, the training details of our CNA model are in Appendix E, and the training details of the autoencoder in Appendix F.
5 Results
Although the figures illustrate individual neuron activity, it is important to clarify that our objective is not to interpret patterns merely as isolated neuron activations, but rather as nets – cohesive, interconnected clusters of active neurons. To emphasize this perspective, Figure 3 illustrates the connectivity among activated neurons in $\boldsymbol{y}^{(S2)}$ (arranged in a circular layout for clarity), after feeding a diagonal line into the network. We select randomly two neurons $n_{i}$ and $n_{j}$ in the left panel and visualize their net fragments $\mathcal{F}(n_{i})$ and $\mathcal{F}(n_{j})$ based on $\boldsymbol{W}^{(L)}$ on the right. This figure highlights that individual neurons (e.g., neurons $n_{i}$ and $n_{j}$) are embedded within broader nets of supportive connections. Without such network-based support, these neurons would be suppressed via attenuation mechanisms and would not appear in the left panel of Figure 3, which displays neuronal activity. Consequently, the subsequent visualizations should be understood as representing supported neurons embedded within nets, even though the underlying connectivity is not explicitly shown.
5.1 Compositionality
Figure 4: The top of the figure illustrates the straight lines used during the training phase. Below, various samples from the evaluation datasets are shown, including kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows $5$-$8$). For each row, two input examples, denoted as $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, are displayed alongside their respective feature activations, $\boldsymbol{y}^{(S1)}$, and neuronal activations, $\boldsymbol{y}^{(S2)}$, shown both without and with addition of Gaussian noise.
The model weights are trained using straight lines (see training data at the top of Figure 4). During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 1-8 of Figure 4). Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows 1 and 2), digits and characters (rows 3 and 4), and line drawings (rows 5–8).
The motivation for why we construct net fragments rather than just using the feature activation $\boldsymbol{y}^{(S1)}$ is demonstrated by introducing additional clutter: For each sample, we generate an input representation both with and without Gaussian noise. This noise corresponds to invalid patterns not observed during training (i.e., are not captured in net fragments), which lack sufficient recurrent support, leading to the deactivation of the corresponding neurons. Thus, the CNA model only allows valid sub-patterns to persist and be composed into novel structures, while invalid patterns, such as noise, are suppressed.
5.2 Filtering Gaussian Noise
Figure 5: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to a lack of support within nets.
To further investigate noise filtering, we introduce Gaussian noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - see Appendix D) to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
The results are visualized in Figure 5: While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated. As additional noise is introduced, more undesired neurons receive support and persist in their activity. However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$. In Appendix G, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
Figure 6: The filtering of (simulated) noise over time: Already at time step $t=0$, most of the noise is suppressed. However, when attenuation increases over time, even more noise is filtered out. The plot is created after adding $10\%$ noise to each feature channel (using a bias $b^{(S2)}=0.7$).
@@ 5.2 Filtering Gaussian Noise
The reason for the filtering is that the additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time. Figure 6 depicts this systematic filtering across timesteps. The figure illustrates a substantial noise reduction already at time step $t=0$ due to initial attenuation. In each subsequent time step, additional clutter is removed. This reduction is due to increased attenuation and falling support as a consequence of the drop-out of supporting neurons.

### v4

@@ 3.2 Stage 2: Net Fragments
Here, the forward connections have the shape $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)},C^{(S1)},h,w}$ and the recurrent connections have the shape $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)},C^{(S2)},h,w}$, where $C^{(S2)}$ is the number of channels in $S2$. Therefore, the pre-synaptic activity $\mathcal{I}$ at time $t$ is increased by the learned amount $\boldsymbol{w}_{j,i}$ from all connected and active neurons ($\boldsymbol{y}_{j,t-1}=1$). Higher support, indicated by larger $\boldsymbol{a}^{\mathcal{I}(S2)}_{i}$, suggests a more likely pattern, as neurons encoding frequent patterns gain reinforcement not only from feedforward connections but also from lateral connections.
Next, we extend this convolutional-based formulation of forward and lateral connections by introducing competitive neurons. Afterward, we describe how the pre-synaptic activity $\boldsymbol{a}^{\mathcal{I}(S2)}$ undergoes attenuation, how it is transformed into $\boldsymbol{y}^{(S2)}$, and how the weight matrices $\boldsymbol{W}^{(F)},\boldsymbol{W}^{(L)}$ are initialized and trained.
Competitive Neurons.
To address the challenge of context entanglement, wherein individual neurons may participate in multiple fragments, we implement competitive neurons as follows: For each base feature channel $c\in C^{(S2)}$, we introduce $\kappa$ parallel alternatives that share the same receptive field structure but are allowed to develop distinct connections. Following our formal definition, each base neuron $n_{c,h^{\prime},w^{\prime}}$ at spatial location $(h^{\prime},w^{\prime})$ has the competitive class:
$\mathcal{C}\left(n_{c,h^{\prime},w^{\prime}}\right)=\left\{n_{c,h^{\prime},w^{\prime}}^{1},\dots,n_{c,h^{\prime},w^{\prime}}^{\kappa}\right\}$ (23)
We implement this by introducing an additional competition dimension of size $\kappa$ to the tensors in $S2$:
$C^{(S2)}$ $\to(C^{(S2)},\,\kappa)$ (24)
$\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)},\,C^{(S1)},\,h,\,w}$ $\to\mathbb{R}^{C^{(S2)},\,\kappa,\,C^{(S1)},\,h,\,w}$
$\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)},\,C^{(S2)},\,h,\,w}$ $\to\mathbb{R}^{C^{(S2)},\,\kappa,\,C^{(S2)},\,h,\,w}$
$\boldsymbol{a}^{\mathcal{I}(L)}\in\mathbb{R}^{C^{(S2)},H,W}$ $\to\mathbb{R}^{C^{(S2)},\,\kappa,\,H,W}$
Although this transformation increases the number of total feature maps by a factor of $\kappa$, only a single competitive neuron per base channel is active at any given spatial location (one neuron within $\mathcal{C}\left(n_{c,h^{\prime},w^{\prime}}\right)$). This sparsity is enforced through a group-wise winner-take-all (WTA) mechanism along the $\kappa$-dimension.
At each time step $t$, competition operates independently within each class $\mathcal{C}\left(n_{c,h^{\prime},w^{\prime}}\right)$. For each base neuron, we compute pre-activation scores for all competitive neurons and select the winning competitive index:
$k^{*}(c,h^{\prime},w^{\prime})\;=\;\operatorname*{arg\,max}_{k\in\{0,\dots,\kappa-1\}}\;\left\langle\boldsymbol{W}_{c,k,:,:,:,:},\;\boldsymbol{z}_{h^{\prime},w^{\prime},t}\right\rangle.$ (25)
where $\boldsymbol{z}_{h^{\prime},w^{\prime},t}$ is the concatenated input patch:
$\boldsymbol{z}_{h^{\prime},w^{\prime},t}\;=\;\begin{bmatrix}\boldsymbol{y}^{(S1)}_{:,\,h^{\prime}:h^{\prime}+H,\,w^{\prime}:w^{\prime}+W}\\ \boldsymbol{y}^{(S2)}_{:,\,h^{\prime}:h^{\prime}+H,\,w^{\prime}:w^{\prime}+W,\,t-1}\end{bmatrix}.$ (26)
Thus, at each spatial location and time step, the index $k^{*}\in\{1,\dots,\kappa\}$ maximizing the correlation between the corresponding filter and the local input patch is chosen. Once the winning index $k^{*}$ is selected, the activation is computed exclusively for the corresponding neuron:
$\boldsymbol{a}^{\mathcal{I}(S2)}_{c,h^{\prime},w^{\prime},t}$ $=\boldsymbol{a}^{\mathcal{I}(F)}_{c,k^{*},h^{\prime},w^{\prime}}+\boldsymbol{a}^{\mathcal{I}(L)}_{c,k^{*},h^{\prime},w^{\prime},t}$ (27)
$\boldsymbol{a}^{\mathcal{I}(F)}_{c,k^{*},h^{\prime},w^{\prime}}$ $=\sum_{c_{\text{in}}=0}^{C^{(S1)}-1}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\boldsymbol{W}^{(F)}_{c,k^{*},\,c_{\text{in}},\,h,\,w}\cdot\boldsymbol{y}^{(S1)}_{c_{\text{in}},\,h^{\prime}+h,\,w^{\prime}+w}$
$\boldsymbol{a}^{\mathcal{I}(L)}_{c,k^{*},h^{\prime},w^{\prime},t}$ $=\sum_{c_{\text{in}}=0}^{C^{(S2)}-1}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\boldsymbol{W}^{(L)}_{c,k^{*},\,c_{\text{in}},\,h,\,w}\cdot\boldsymbol{y}^{(S2)}_{c_{\text{in}},\,h^{\prime}+h,\,w^{\prime}+w,\,t-1}$
Initially, all competitive neurons $n_{c,h^{\prime},w^{\prime}}^{k}$ are identical due to shared initialization. However, Hebbian learning induces specialization: each competitive neuron within $\mathcal{C}\left(n_{c,h^{\prime},w^{\prime}}\right)$ gradually diverges, learning to encode different contextual variations of the same base feature.
In a practical implementation, we merge the $\kappa$ and channel dimensions, resulting in a flat channel axis of size $C^{(S2)}\cdot\kappa$. This allows the use of efficient standard 3D convolutional layers, where non-winning neurons are simply masked out at each spatial location by setting them to 0, preserving the theoretical definition.
@@
The pre-synaptic activity is processed by an activation function $\mathcal{A}\left(\boldsymbol{a}^{\mathcal{I}},\mathcal{E}\right)$. In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{I}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range [0,1][0,1][ 0 , 1 ] (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active. Specifically, attenuation is implemented as:
@@
Weight Initialization.
The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are initially copied into all competitive $S2$ neurons in the same position. This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to 111 if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of competitive neurons. The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to 111.
Learning of Weights.
Support within net fragments is learned via Hebbian plasticity, so that frequently occurring patterns that lead to correlated activation patterns between neurons $n_{i}\sim n_{j}$ are “stored” in the system and receive more support:
Here, $\eta$ represents the learning rate and $\boldsymbol{W}^{(S2)}$ stands for both $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$. Note that the same convolutional kernel is applied at various spatial positions. The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects (across all spatial positions). If the resulting average is positive, the corresponding connection increases as it connects more simultaneously active neurons than neurons that fire disjointly. Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced. Importantly, the weight update is applied only once per sample, after the full sequence of $T$ timesteps has been processed. Learning proceeds over multiple epochs, meaning each sample is revisited multiple times. Due to the small learning rate of $\eta=0.2$, these weight changes accumulate gradually.
4 Experimental Results
4.1 Experimental Setup
Our experiments are designed to rigorously probe the core representational capability of the CNA: its ability to form net fragments during training and recombine them into larger nets during inference. To isolate this property, we deliberately avoid task-specific or dataset-driven complexity and instead focus on the model’s behavior in its most fundamental setting.
We therefore train CNA on binary images of straight lines at varying orientations. Straight lines serve as the most elementary visual primitives: they are simple enough to avoid confounds yet naturally compose into richer structures such as line drawings of objects. This controlled regime directly tests whether CNA can learn fragmentary representations and later reassemble them into unseen global patterns. By contrast, benchmarks such as MNIST Lecun et al. (1998), CIFAR Krizhevsky and Hinton (2009), or ImageNet Deng et al. (2009) involve dense, texture-heavy images where the link between local features and global structure is opaque. Evaluating CNA on such data at this stage would obscure its core representational mechanism, and, as we discuss in Section 6, would require introducing an additional object-representing area and more advanced filters, which we defer to future work.
Evaluation targets two key properties:
  • Compositionality: Can learned line fragments recombine into novel structures, such as object line drawings unseen during training?
  • Robustness: Can nets be retrieved reliably under noise or occlusion?
To this end, we present the model after training and without updating any parameters with (i) line drawings of objects composed of straight-line segments, (ii) inputs corrupted with Gaussian noise, and (iii) partially occluded patterns.
For comparison, we also train a conventional autoencoder on the same line-stimuli dataset to highlight representational differences. Further experimental details are provided in the appendix: We describe the training and testing datasets in more detail in Appendix D, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ in Appendix A, the training details of our CNA model in Appendix E, and the training details of the autoencoder in Appendix F.
4.2 Interpreting Figures: General Considerations
A defining feature of CNA is that it represents patterns not through individual neuron activations, but through nets – cohesive, interconnected clusters of mutually supportive neurons. This network-based organization means that an individual neuron’s activation is meaningful only in the context of its embedding within a larger net. Understanding CNAs’ outputs, therefore, requires a shift in perspective: from isolated neuron activity to the connectivity and structure of the nets they form.
Figure 3 illustrates this principle. After presenting a diagonal line input, the left panel shows the resulting neuronal activity $\boldsymbol{y}^{(S2)}$, while the right panel visualizes the underlying net structure derived from recurrent connections $\boldsymbol{W}^{(L)}$. For clarity, neurons are arranged in a circular layout, and the net fragments $\mathcal{F}(n_{i})$ and $\mathcal{F}(n_{j})$ of two randomly selected neurons $n_{i}$ and $n_{j}$ are highlighted in color.
This visualization demonstrates two important points. First, individual neurons (e.g., $n_{i}$ and $n_{j}$) remain active only if embedded within supportive nets: without sufficient recurrent support, they are suppressed by attenuation mechanisms and would appear inactive (black) in the left panel. Second, this shows that observed activity patterns should always be interpreted as the visible footprint of underlying nets, even when those nets are not explicitly drawn.
Consequently, all subsequent visualizations of neural activity must be understood in this net-based context: the squares represent neurons that survive attenuation precisely because they are supported by their respective nets.
All our visualizations of neural activity are color-coded according to the competing channel groups $\mathcal{C}(n_{i})$. Each color corresponds to a feature type (e.g., horizontal, vertical, or diagonal lines; see Appendix A). However, these colors do not distinguish between different competitive channels $k\in{1,\dots,\kappa}$ within the same feature group, meaning that two neurons with the same color may belong to different competitive subnetworks with distinct connectivity patterns.
4.3 Compositionality
Figure 4: Training and evaluation data. The top of the figure illustrates the straight lines used during the training phase. Below, various samples from the evaluation datasets are shown, including kinked lines (rows 111 and 222), digits and characters (rows 333 and 444), and line drawings (rows 555888). For each row, two input examples, denoted as $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, are displayed alongside their respective feature activations, $\boldsymbol{y}^{(S1)}$, and neuronal activations, $\boldsymbol{y}^{(S2)}$, shown both without and with addition of Gaussian noise.
The model weights are trained using straight lines (see training data examples at the top of Figure 4). During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 111888 of Figure 4). Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows 111 and 222), digits and characters (rows 333 and 444), and line drawings (rows 555888).
4.4 Filtering Gaussian Noise
The motivation for why we construct net fragments rather than just using the feature activation $\boldsymbol{y}^{(S1)}$ is demonstrated by introducing additional clutter: For each sample, we generate an input representation both with and without Gaussian noise. This noise corresponds to invalid patterns not observed during training (i.e., are not captured in net fragments) that consequently lack sufficient recurrent support, leading to the deactivation of the corresponding neurons. Thus, the CNA model only allows valid sub-patterns to persist and be composed into novel structures, while invalid patterns, such as noise, are suppressed.
Figure 5: Noise filtering efficacy: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to a lack of support within nets.
@@ 4.4 Filtering Gaussian Noise
[…]
@@
The reason that the filtering works is that the additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time. Figure 6 depicts this systematic filtering across timesteps. The figure illustrates a substantial noise reduction already at time step $t=0$ due to initial attenuation. In each subsequent time step, additional clutter is removed. This reduction is due to increased attenuation and falling support as a consequence of the drop-out of supporting neurons.